A 4.00 kg mass is connected to a spring with a sprshow a solution please
Answer #1

angular frequency = sqrt(k / m)

angular frequency = sqrt(9 / 4)

angular frequency = 1.5 rad/sec

x(t) = 16 * cos(wt)

x(t) = amplitude * cos(wt)
so,

Amplitude = 16 cm

angular frequency = 2 * pi * frequency

1.5 = 2 * pi * frequency

frequency = 0.2387 Hz

time period = 1 / frequency

time period = 1 / 0.2387

time period = 4.189 sec

maximum velocity = w * amplitude

maximum velocity = 1.5 * 0.16

maximum velocity = 0.24 m/s

energy = 0.5 * k * amplitude^2

energy = 0.5 * 9 * 0.16^2

energy = 0.1152 J